OpenStudy (anonymous):

Graph the function of f f(x)=-6x^2 ;x greater than/equal to 0

OpenStudy (poopsiedoodle):

try plugging it in here: http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html if it's greater than or = to 0, then you shade above the line, and it's solid.

OpenStudy (mathmale):

What aspect of this graphing assignment is causing you trouble? Would you be able to graph y=x^2 first? y=6x^2? y=-6x^2? Working through these in that order may give you a better understanding of what to do.

OpenStudy (anonymous):

is it graphed as a regular parabola?

OpenStudy (mathmale):

Yes, absolutely, except that y ou'd focus only on the part of the graph that corresponds to x=0 or x>0. y=x^2 is a "regular parabola"

OpenStudy (anonymous):

Okay thanks